The college of business at Kerouac University is planning to begin an online MBA program. The
initial start-up cost for computing equipment, facilities, course development and staff recruitment
and development is $400,000. The college plans to charge tuition of $20,000 per student per year.
However, the university administration will charge the college $10,000 per student for the first 100
students enrolled each year for administrative costs and its share of the tuition payments.
a. How many students does the college need to enroll in the first year to break-even?
b. If the college can enroll 80 students the first year, how much profit will it make?
MAT540 Homework
Week 1
Page 2 of 3
c. The college believes it can increase tuition to $25,000, but doing so would reduce enrollment to
50. Should the college consider doing this?
1
Expert's answer
2015-10-16T00:01:10-0400
The initial start-up cost is $400,000, tuition of $20,000 per student per year. However, the university administration will charge the college $10,000 per student for the first 100 students enrolled each year for administrative costs and its share of the tuition payments.
a. The college needs to enroll in the first year (400,000+10,000*100)/20,000 = 70 students to break-even.
b. If the college can enroll 80 students the first year, its profit will be: TP = 20,000*80 - (400,000+10,000*100) = $200,000
c. If the college believes it can increase tuition to $25,000, but doing so would reduce enrollment to 50, then profit will be TP = 25,000*50 - (400,000+10,000*100) = -$150,000, so the college shouldn't consider doing this.